Problem 4 The Economist magazine uses the price of Big Macs in different countries to calculate PPP (or LOOP) exchange rates. It is called the Big Mac Index. The following data are from the January 2018 copy of the Economist. I Big Mac Price (local currency) USD 5.28 Country United States Australia New Zealan AUD 5.90 NZD 6.2 Calculate the implied LOOP values for the US-Australian dollar exchange rate and the NZ-Australian dollar exchange rate. . Answer:
On January 30, 2018 the respective market values for the USD/AUD and NZD/AUD exchange rates were: USD/AUD 0.808 and NZD/AUD = 1.093 Relative to the two LOOP values was the Australian dollar over or under-valued in January 2018? Calculate the degree of over or under-valuation in percent. 2. Answer
Answer #1

4.

Implied US-AUD exchange rate = 5.28/5.90

Implied US-AUD exchange rate = .895

Implied NZ-AUD exchange rate = 6.2/5.90

Implied NZ-AUD exchange rate = 1.051

2.

Against US dollar, AUD is undervalued as it is weaker in market, but implied value is strong.

Degree of undervaluation = (.808/895 -1) = 9.72% (-ve)

Against NZ dollar, AYD is overvalued as it is stronger in market, but implied value is weak.

Degree over overvaluation = 1.093/1.051 – 1 = 4%